# SynTable: A Synthetic Data Generation Pipeline for Unseen Object Amodal Instance Segmentation of Cluttered Tabletop Scenes

### v2

@@ 4.1 Object Models Used in Generating SynTable-Sim
We use 1075 object CAD models from the Google Scanned Objects dataset [9] and the Benchmark for 6D Object Pose Estimation (BOP) [14] for generating our train dataset. The Google Scanned Objects dataset features over 1030 photorealistic 3D scanned household objects with real-life dimensions, and BOP features 3D object models from household and industrial objects. As the context of our work is in robotic grasping, we only select objects that a parallel gripper can grasp to be part of our dataset. Upon inspection of the Google Scanned Objects dataset, we filter out invalid objects that contain more than two instances in each model and keep the remaining 891 valid objects for our training dataset. From the BOP, we exclude 21 objects from the YCB-Video dataset that we include in our validation dataset and use the remaining 184 objects for our training dataset. We also create a synthetic validation set using 78 novel objects from the YCB dataset [4]. We sample a table object from 10 Omniverse Nucleus table assets to provide randomization for each scene. To load the 3D object models into Isaac Sim, we convert the OBJ and texture files to the Universal Scene Description (USD) format.
@@ 4.2 Dataset Configuration
With 50 viewpoints for each scene, we generated 900 scenes to create 45,000 RGB-D images for the training dataset and 100 scenes to create 5,000 RGB-D images for the validation dataset. $N_{lower}=1$ to $N_{upper}=40$ objects are rendered on randomly textured tabletop planes in each scene. We used 130 materials from the Omniverse Nucleus material assets to be applied randomly on the walls, floor, and table for domain randomization purposes. $L_{lower}=0$ to $L_{upper}=2$ spherical lights are sampled for each scene. The viewpoint and lighting hemisphere parameters are automatically sampled based on the table dimensions. The camera parameters used are horizontal aperture: 2.63, vertical aperture: 1.96, and focal length: 1.88 to mimic the configuration of the RealSense LiDAR Camera L515. The rest of the parameters follow the default configurations of the pipeline.
@@ 4.3 Syntable-Sim Versus Other Cluttered Tabletop Datasets
We compare our SynTable-Sim dataset with several other cluttered tabletop datasets from previous works in Table 1. Our tabletop dataset is the only one providing all annotations related to amodal instance segmentation. Our dataset also has the most extensive variety of objects, the highest number of occlusion instances, and the second highest average occlusion rate. These characteristics make our dataset very challenging for amodal instance segmentation tasks.
@@ 5 Experiments
In this section, we present the results of our experiments aimed at evaluating the effectiveness of our dataset generation pipeline in producing synthetic datasets with good Sim-to-Real transfer performance. We use our SynTable-Sim sample dataset to train a state-of-the-art (SOTA) UOAIS model, UOAIS-Net [2]. UOAIS-Net is then evaluated on the SynTable-Sim validation set and OSD-Amodal [2] test set. To verify to consistency of our results and also to further demonstrate the capability of SynTable to improve the performance of a variety of different UOAIS models, we also train and evaluate three other UOAIS models—Amodal MRCNN [12], ORCNN [12], ASN [21]—on the SynTable-Sim and OSD-Amodal datasets respectively. The results and analyses for these models are included in Section 12 of our supplementary materials.
@@ 5.1 Training Strategy
We train UOAIS-Net on the UOAIS-Sim tabletop and SynTable-Sim datasets using an NVIDIA Tesla V100 GPU with 16 GB of memory. For both datasets, we use 90% of the images for training and 10% for validation. To train UOAIS-Net using the UOAIS-Sim tabletop dataset, we use the same hyperparameters as Back et al. [2]. To train UOAIS-Net with SynTable-Sim, we modified the depth range hyperparameter, which is used to preprocess input depth images. Specifically, we changed the range from the 2500 mm to 40000 mm range set by Back et al. to a narrower range of 250 mm to 2500 mm. This adjustment is required because our dataset reflects real-world proportions and has a smaller depth range than the UOAIS-Sim dataset. We also use a similar training strategy for training Amodal MRCNN, ORCNN, and ASN.
@@ 5.2 Evaluation Metrics
[…]
Due to the subjectivity of the objects’ invisible masks, evaluating UOAIS model performance solely based on the overlap and boundary P/R/F of segmented objects may be inaccurate. The current UOAIS occlusion evaluation metrics measure how well the model can predict if individual objects are occluded. However, these metrics do not measure object occlusion ordering in a scene which would be helpful in scene understanding for robotic grasp planning. The OOAM can represent a scene’s occlusion order, while the OODG can be derived from the OOAM. By knowing the occlusion order in a scene, a robot can use topological sorting of the predicted OODG to plan the order of grasping to reach the occluded object of interest. Thus, to evaluate the accuracy of object occlusion order in an image, we introduce a new metric called Occlusion Order Accuracy ($ACC_{OO}$) as stated in Equation 8.
$\small ACC_{OO}=\frac{sum(similarityMatrix)-gtOOAMDiagonalSize}{gtOOAMSize-gtOOAMDiagonalSize}$ (8)
In Equation 8, $similarityMatrix$ is the element-wise equality comparison between the ground truth OOAM, $gtOOAM$, and the predicted OOAM, $predOOAM$. As an object cannot occlude itself, the diagonal of any OOAM is always 0. Thus, we subtract the number of elements along the diagonal of $gtOOAM$, $gtOOAMDiagonalSize$, from the calculation of $ACC_{OO}$. $ACC_{OO}$ is used to evaluate the model’s ability to accurately determine the order of occlusions in a clutter of objects by comparing the OOAM generated by the model to the ground truth OOAM using Algorithm 2. We give a specific example of how to compute $ACC_{OO}$ in Sections 9 and 10 of our supplementary materials.
@@ 5.3 Results
CR HF CA
PD
[…]
OV BO F@.75
1 UOAIS-Sim (Tabletop)
42.4
34.1
47.1
21.6
15.2
18.5
43.1
61.8
42.5
32.3
37.1
12.7
1 SynTable-Sim (Ours) 80.9
61.8
78.1 52.4 31.2
41.3
75.7 86.7 81.1
64.3
74.4 82.9
1 SynTable-Sim-0.5X (Ours)
80.7
63.8
77.3
51.9
30.2
42.9 75.7
84.1
80.5
65.4
71.7
82.7
2 UOAIS-Sim (Tabletop)
0.8
26.1
33.1
66.7
15.5
7.7
20.4
60.8
78.1
25.9
27.6
51.8
42.7
2 SynTable-Sim (Ours)
0.8
67.7
56.0
81.2
49.4
30.1
48.6
72.5
89.8
71.8
61.3
78.2
86.6
2 SynTable-Sim-0.5X (Ours)
0.8
75.6 61.2 83.5 53.6 31.2
48.5
75.5 90.1 76.8 64.5 78.3 87.0
3 UOAIS-Sim (Tabletop)
0.8
71.8
62.8
81.4
55.6 31.3 44.6 75.1
86.2
70.2
63.2
73.2
79.6
3 SynTable-Sim (Ours)
0.8
78.3
58.8
81.9
54.0
29.7
43.9
66.6
93.2
79.2
60.4
77.2
87.7
3 SynTable-Sim-0.5X (Ours)
0.8
74.0
57.5
83.3
49.2
23.9
41.0
65.7
93.4
74.2
59.2
79.2
87.6
4 UOAIS-Sim (Tabletop)
0.5
49.0
50.3
82.7
42.3
23.9
40.3
68.9
84.0
47.3
50.0
70.6
80.4
4 SynTable-Sim (Ours)
0.5
64.4 51.5
84.3
47.3 24.2
47.4
60.0
91.9 65.3 53.7 78.2
87.0
4 SynTable-Sim-0.5X (Ours)
0.5
55.0
47.2
85.9
43.2
22.0
48.4
55.4
91.5
55.3
46.6
76.9
87.8
@@ 4 SynTable-Sim (Ours)
[…]
P R
F
P R
F
P R
F
P R
F
P R
F
P R
F
UOAIS-Sim
35.9 65.4 42.4 31.4 42.8 34.1 47.1 55.9 24.5 21.6 45.3 19.3 15.2 18.5 36.2 61.3 42.5 30.8 39.2 32.3 37.1 43.1 61.8 12.7
@@
[…]
OV BO F@.75
UOAIS-Sim (Tabletop)
38.0
37.8
35.9
14.1
12.9
7.6
47.2
72.9
40.4
38.9
34.8
31.6
@@
Table 2 compares the performance of UOAIS-Net on the OSD-Amodal dataset after training on the UOAIS-Sim tabletop dataset and our SynTable-Sim sample dataset. We conduct four sets of experiments. In each set of experiments, we vary the amount of data augmentation used and the size of the dataset we use for training. In our first set of experiments, we can see that the UOAIS-Net trained on the SynTable-Sim dataset significantly outperforms the UOAIS-Net trained on the UOAIS-Sim tabletop dataset in all metrics. Even when we train UOAIS-Net using a dataset of the same size as UOAIS-Sim (SynTable-Sim-0.5X), the performance is still remarkably better than the UOAIS-Net trained on the UOAIS-Sim tabletop dataset across all metrics. A detailed breakdown of the precision P, recall R, and F-measure F, and F@.75 scores for the amodal, invisible, and visible masks for our first set of experiments are shown in Table 3. We observe that except for the Boundary precision scores of the invisible masks, UOAIS-Net achieves substantial improvements in all other metrics.
@@
Similarly, from Table 4, the UOAIS-Net model trained on the SynTable-Sim dataset also outperforms the one trained on UOAIS-Sim tabletop dataset in all metrics when both models are benchmarked on SynTable-Sim validation dataset.
Our experiments demonstrate the effectiveness of our proposed dataset generation pipeline, SynTable, in improving the Sim-to-Real transfer performance of SOTA deep learning computer vision models for UOAIS. These results also highlight the potential of SynTable for addressing the challenge of annotating amodal instance segmentation masks.

### v3

@@ 4.1 Object Models Used in Generating SynTable-Sim
We use 1075 object CAD models from the Google Scanned Objects dataset [8] and the Benchmark for the 6D Object Pose Estimation (BOP) [13] to generate our train dataset. The Google Scanned Objects dataset features more than 1030 photorealistic 3D scanned household objects with real-life dimensions, and BOP features 3D object models from household and industrial objects. Upon inspection of the Google Scanned Objects dataset, we filter out invalid objects that contain more than two instances in each model and keep the remaining 891 valid objects for our training dataset. From the BOP, we exclude 21 objects from the YCB-Video dataset that we include in our validation dataset and use the remaining 184 objects for our training dataset. We also create a synthetic validation set using 78 novel objects from the YCB dataset [4]. We sample a table object from 10 Omniverse Nucleus table assets to provide randomization for each scene. To load the 3D object models into Isaac Sim, we converted the OBJ and texture files to the Universal Scene Description (USD) format.
@@ 4.2 Dataset Configuration
With 50 viewpoints for each scene, we generated 900 scenes to create 45,000 RGB-D images for the training dataset and 100 scenes to create 5,000 RGB-D images for the validation dataset. $N_{lower}=1$ to $N_{upper}=40$ objects are rendered in randomly textured tabletop planes in each scene. We used 130 materials from Omniverse Nucleus material assets to be randomly applied on the walls, floor, and table for domain randomization purposes. $L_{lower}=0$ to $L_{upper}=2$ spherical lights are sampled for each scene. The viewpoint and lighting hemisphere parameters are automatically sampled based on the table dimensions. The camera parameters used are horizontal aperture: 2.63, vertical aperture: 1.96, and focal length: 1.88 to mimic the configuration of the RealSense LiDAR Camera L515. The rest of the parameters follow the default configurations of the pipeline.
@@ 4.3 Syntable-Sim Versus Other Cluttered Tabletop Datasets
We compare our SynTable-Sim dataset with several existing cluttered tabletop datasets in Table 1. Our tabletop dataset is the only one that provides complete annotations for all aspects of amodal instance segmentation. Furthermore, our dataset contains the most extensive variety of objects, the highest number of occlusion instances, and the second highest average occlusion rate — critical factors that significantly enhance the complexity and realism of training scenarios. These characteristics make our dataset very challenging for amodal instance segmentation tasks.
Additionally, SynTable-Sim exhibits a significantly higher proportion of heavily occluded objects in its training set compared to UOAIS-Sim, aligning more closely with the OSD-Amodal dataset, as shown in Figure 7 in the supplementary materials. This high occlusion density ensures that models trained on our dataset generalize better to real-world cluttered environments. Moreover, the weakly connected component size, which quantifies the number of mutually overlapped regions per OODG and serves as a metric for scene complexity [15], is consistently larger in SynTable-Sim compared to UOAIS-Sim (Figure 8 in the supplementary materials). This indicates that our dataset presents significantly more intricate occlusion patterns, enabling amodal segmentation models to learn more robust occlusion reasoning capabilities.
@@ 5 Experiments
In this section, we present the results of our experiments aimed at evaluating the effectiveness of our dataset generation pipeline in producing synthetic datasets with good Sim-to-Real transfer performance. We used our SynTable-Sim sample dataset to train a state-of-the-art (SOTA) UOAIS model, UOAIS-Net [2]. UOAIS-Net is evaluated on the SynTable-Sim validation set and the OSD-Amodal [2] test set. To verify consistency of our results and further demonstrate the capability of SynTable to improve the performance of a variety of different UOAIS models, we also train and evaluate three other UOAIS models—Amodal MRCNN [11], ORCNN [11], ASN [21]—on the SynTable-Sim and OSD-Amodal datasets respectively.
@@ 5.1 Training Strategy
We train UOAIS-Net on the UOAIS-Sim tabletop and SynTable-Sim datasets using an NVIDIA Tesla V100 GPU with 16 GB of memory. For both datasets, we used 90% of the images for training and 10% for validation. To train UOAIS-Net using the UOAIS-Sim tabletop dataset, we use the same hyperparameters as Back et al. [2]. To train UOAIS-Net with SynTable-Sim, we modified the depth range hyperparameter, which is used to preprocess input depth images. Specifically, we changed the range from the 2500 mm to 40000 mm range set by Back et al. to a narrower range of 250 mm to 2500 mm. This adjustment is required because our dataset reflects real-world proportions and has a smaller depth range than the UOAIS-Sim dataset. We also use a similar training strategy to train Amodal MRCNN, ORCNN, and ASN.
@@ 5.2 Evaluation Metrics
We measure the performance of UOAIS-Net on the following traditional metrics [29, 5, 19]: Overlap P/R/F, Boundary P/R/F, and F@.75 for the amodal, visible, and invisible masks. Overlap P/R/F and Boundary P/R/F evaluate the whole area and the sharpness of the prediction, respectively, where P, R, and F are the precision, recall, and F-measure of instance masks after the Hungarian matching, respectively. F@.75 is the percentage of segmented objects with an Overlap F-measure greater than 0.75. We also report the accuracy ($\textit{ACC}_{\mathcal{O}}$) and F-measure ($\textit{F}_{\mathcal{O}}$) of occlusion classification, where $\textit{ACC}_{\mathcal{O}}=\frac{\delta}{\alpha}$, $\textit{F}_{\mathcal{O}}=\frac{2P_{o}R_{o}}{P_{o}+R_{o}}$, $P_{o}=\frac{\delta}{\beta}$, $R_{o}=\frac{\delta}{\gamma}$. $\alpha$ is the number of the matched instances after the Hungarian matching. $\beta$, $\gamma$, and $\delta$ are the number of occlusion predictions, ground truths, and correct predictions, respectively. We provide more details about the evaluation metrics in Section 9 of our supplementary materials.
Due to the subjectivity of the invisible masks of objects, the evaluation of the performance of the UOAIS model solely based on the overlap and boundary P/R/F of segmented objects may be inaccurate. The current UOAIS occlusion evaluation metrics measure how well the model can predict whether individual objects are occluded. However, these metrics neglect hierarchical occlusion relationships, which are crucial for systems requiring structured scene understanding. The Occlusion Order Adjacency Matrix (OOAM) encodes these relationships, and the derived Occlusion Order Directed Graph (OODG) enables applications such as sequencing interactions in cluttered environments (for example, retrieving obscured items) or rendering occluded objects in augmented reality. To quantify a model’s ability to infer occlusion hierarchies, we propose the Occlusion Order Accuracy Occlusion Order Accuracy ($ACC_{OO}$) metric as defined in Equation 1.
$ACC_{OO}=\frac{sum(similarityMatrix)-gtOOAMDiagonalSize}{gtOOAMSize-gtOOAMDiagonalSize}$ (1)
In Equation 1, $similarityMatrix$ is the element-wise equality comparison between the ground truth OOAM, $gtOOAM$, and the predicted OOAM, $predOOAM$. As an object cannot occlude itself, the diagonal of any OOAM is always 0. Thus, we subtract the number of elements along the diagonal of $gtOOAM$, $gtOOAMDiagonalSize$, from the calculation of $ACC_{OO}$. $ACC_{OO}$ is used to evaluate the model’s ability to accurately determine the order of occlusions in a clutter of objects by comparing the OOAM generated by the model to the ground truth OOAM using Algorithm 2. We give a specific example of how to compute $ACC_{OO}$ in Sections 10 and 11 of our supplementary materials.
@@ 5.3 Results
CR HF CA PD OV BO F@.75 OV BO F@.75 $\textit{F}_{\mathcal{O}}$ $\textit{ACC}_{\mathcal{O}}$ OV BO F@.75
1 UOAIS-Sim (Tabletop) 42.4 34.1 47.1 21.6 15.2 18.5 43.1 61.8 42.5 32.3 37.1 12.7
1 SynTable-Sim (Ours) 80.9 61.8 78.1 52.4 31.2 41.3 75.7 86.7 81.1 64.3 74.4 82.9
1 SynTable-Sim-0.5X (Ours) 80.7 63.8 77.3 51.9 30.2 42.9 75.7 84.1 80.5 65.4 71.7 82.7
2 UOAIS-Sim (Tabletop) 0.8 26.1 33.1 66.7 15.5 7.7 20.4 60.8 78.1 25.9 27.6 51.8 42.7
2 SynTable-Sim (Ours) 0.8 67.7 56.0 81.2 49.4 30.1 48.6 72.5 89.8 71.8 61.3 78.2 86.6
2 SynTable-Sim-0.5X (Ours) 0.8 75.6 61.2 83.5 53.6 31.2 48.5 75.5 90.1 76.8 64.5 78.3 87.0
3 UOAIS-Sim (Tabletop) 0.8 71.8 62.8 81.4 55.6 31.3 44.6 75.1 86.2 70.2 63.2 73.2 79.6
3 SynTable-Sim (Ours) 0.8 78.3 58.8 81.9 54.0 29.7 43.9 66.6 93.2 79.2 60.4 77.2 87.7
3 SynTable-Sim-0.5X (Ours) 0.8 74.0 57.5 83.3 49.2 23.9 41.0 65.7 93.4 74.2 59.2 79.2 87.6
4 UOAIS-Sim (Tabletop) 0.5 49.0 50.3 82.7 42.3 23.9 40.3 68.9 84.0 47.3 50.0 70.6 80.4
4 SynTable-Sim (Ours) 0.5 64.4 51.5 84.3 47.3 24.2 47.4 60.0 91.9 65.3 53.7 78.2 87.0
4 SynTable-Sim-0.5X (Ours) 0.5 55.0 47.2 85.9 43.2 22.0 48.4 55.4 91.5 55.3 46.6 76.9 87.8
@@
Overlap Boundary F@.75 Overlap Boundary F@.75 Overlap Boundary F@.75 $\textit{F}_{\mathcal{O}}$ $\textit{ACC}_{\mathcal{O}}$
P R F P R F P R F P R F P R F P R F
UOAIS-Sim 35.9 65.4 42.4 31.4 42.8 34.1 47.1 55.9 24.5 21.6 45.3 19.3 15.2 18.5 36.2 61.3 42.5 30.8 39.2 32.3 37.1 43.1 61.8 12.7
@@
OV BO F@.75 OV BO F@.75 $\textit{F}_{\mathcal{O}}$ $\textit{ACC}_{\mathcal{O}}$ OV BO F@.75
UOAIS-Sim (Tabletop) 38.0 37.8 35.9 14.1 12.9 7.6 47.2 72.9 40.4 38.9 34.8 31.6
@@
Table 2 compares the performance of UOAIS-Net on the OSD-Amodal dataset after training on the UOAIS-Sim tabletop dataset and our SynTable-Sim sample dataset. We conducted four sets of experiments. In each set of experiments, we vary the amount of data augmentation used and the size of the dataset we use for training.
In our first set of experiments, we can see that the UOAIS-Net trained on the SynTable-Sim dataset significantly outperforms the UOAIS-Net trained on the UOAIS-Sim tabletop dataset in all metrics. Even when we train UOAIS-Net using a dataset of the same size as UOAIS-Sim (SynTable-Sim-0.5X), the performance is still remarkably better than the UOAIS-Net trained on the UOAIS-Sim tabletop dataset across all metrics. A detailed breakdown of the precision P, recall R, and F-measure F, and F@.75 scores for the amodal, invisible and visible masks for our first set of experiments is shown in Table 3. We observe that except for the Boundary precision scores of the invisible masks, UOAIS-Net achieves substantial improvements in all other metrics.
@@
Similarly, from Table 4, the UOAIS-Net model trained on the SynTable-Sim dataset outperforms the one trained on UOAIS-Sim tabletop dataset in all metrics when both models are benchmarked on SynTable-Sim validation dataset.
We evaluated the effectiveness of SynTable-Sim across different UOAIS models comprising distinct architectures. Table 5 compares the performance of UOAIS models— Amodal MRCNN, ORCNN, ASN, and UOAIS-Net—on the OSD-Amodal dataset after training on the UOAIS-Sim tabletop dataset and our SynTable-Sim sample dataset. For each model result in our experiments, we used seed 7 for training. Generally, across most metrics, the UOAIS models trained on SynTable-Sim outperform the same models trained on the UOAIS-Sim tabletop dataset. There is also a significant improvement in the results of $ACC_{oo}$ for Amodal MRCNN, ORCNN, and ASN when trained on our SynTable-Sim as compared to the UOAIS-Sim tabletop dataset. This is consistent with the performance trend observed for UOAIS-Net and, therefore, demonstrates that SynTable is an effective tool for generating high-quality datasets that can improve the performance of UOAIS models. A detailed breakdown of the precision P, recall R, and F-measure F, and F@.75 scores for the amodal, invisible, and visible masks are shown in Table 6.
OV BO F@.75 OV BO F@.75 $\textit{F}_{\mathcal{O}}$ $\textit{ACC}_{\mathcal{O}}$ OV BO F@.75
UOAIS-Sim (Tabletop) Amodal MRCNN 36.7 26.9 45.7 8.8 4.8 7.7 39.2 54.8 38.7 26.3 32.2 15.6
ORCNN 36.3 25.4 47.0 12.2 6.7 9.0 43.8 59.2 30.5 21.8 29.6 21.5
ASN 40.5 33.6 49.8 17.4 12.1 15.0 47.0 63.2 39.3 31.6 36.8 17.8
UOAIS-Net 49.0 50.3 82.7 42.3 23.9 40.3 68.9 84.0 47.3 50.0 70.6 80.4
SynTable-Sim (Ours) Amodal MRCNN 74.5 57.5 77.2 41.3 23.5 37.6 69.3 79.4 73.8 57.7 66.1 79.2
ORCNN 74.2 58.2 77.1 44.7 24.3 33.8 72.9 82.2 72.0 58.3 67.7 79.1
ASN 78.2 60.2 75.3 46.4 27.7 35.8 72.6 83.0 78.1 61.8 68.9 80.2
UOAIS-Net 64.4 51.5 84.3 47.3 24.2 47.4 60.0 91.9 65.3 53.7 78.2 87.0
Training Set Method Amodal Mask Invisible Mask Visible Mask
Overlap Boundary F@.75 Overlap Boundary F@.75 Overlap Boundary F@.75
P R F P R F P R F P R F P R F P R F
UOAIS-Sim (Tabletop) Amodal 27.9 66.7 36.7 22.5 39.8 26.9 45.7 20.2 24.9 8.8 16.4 19.9 4.8 7.7 30.1 60.5 38.7 22.0 37.8 26.3 32.2
MRCNN
ORCNN 26.3 71.1 36.3 19.8 42.4 25.4 47.0 41.5 22.7 12.2 33.9 17.9 6.7 9.0 21.4 63.4 30.5 16.6 38.2 21.8 29.6
ASN 31.7 67.8 40.5 28.6 45.4 33.6 49.8 47.6 23.4 17.4 38.8 20.2 12.1 15.0 32.0 63.5 39.3 28.2 41.5 31.6 36.8
UOAIS-Net 37.2 85.5 49.0 41.1 71.3 50.3 82.7 50.9 54.0 42.3 24.8 41.1 23.9 40.3 35.4 81.6 47.3 41.3 69.3 50.0 70.6
SynTable-Sim (Ours) Amodal 72.3 81.6 74.5 54.6 64.5 57.5 77.2 54.9 48.0 41.3 30.3 38.8 23.5 37.6 72.1 78.4 73.8 55.1 63.7 57.7 66.1
MRCNN
ORCNN 73.7 80.8 74.2 55.6 64.5 58.2 77.1 61.0 47.1 44.7 31.1 38.6 24.3 33.8 69.8 79.1 72.0 55.7 64.3 58.3 67.7
ASN 78.2 80.3 78.2 57.8 64.9 60.2 75.3 65.2 46.2 46.4 32.9 38.7 27.7 35.8 77.5 80.1 78.1 60.4 65.4 61.8 68.9
UOAIS-Net 53.9 86.3 64.4 40.9 74.6 51.5 84.3 53.0 60.0 47.3 20.5 48.3 24.2 47.4 55.0 86.2 65.3 43.2 75.3 53.7 78.2
OV BO F@.75 OV BO F@.75 $\textit{F}_{\mathcal{O}}$ $\textit{ACC}_{\mathcal{O}}$ OV BO F@.75
UOAIS-Sim (Tabletop) Amodal MRCNN 27.1 25.2 23.8 6.7 6.2 3.5 35.8 66.0 29.1 26.2 23.5 19.0
ORCNN 30.9 29.0 28.1 12.5 11.4 8.0 39.9 68.4 31.8 30.2 27.3 23.1
ASN 33.3 34.4 35.3 10.3 9.1 5.0 47.6 72.3 35.0 36.0 34.1 31.6
UOAIS-Net 39.9 40.5 38.6 17.0 15.5 9.6 49.6 74.9 41.6 40.7 35.9 31.6
SynTable-Sim (Ours) Amodal MRCNN 83.5 76.2 72.5 35.4 31.8 16.4 73.2 80.3 85.7 79.1 71.1 72.8
ORCNN 83.4 76.0 72.2 34.4 29.3 15.3 67.2 73.7 85.3 78.9 70.9 73.0
ASN 83.6 76.9 73.9 38.5 35.1 18.5 74.8 81.5 86.1 80.0 72.8 75.8
UOAIS-Net 83.7 77.5 75.1 40.3 36.7 20.2 75.5 82.0 86.2 80.1 73.3 77.4
Our experiments demonstrate the effectiveness of our proposed dataset generation pipeline, SynTable, in improving the Sim-to-Real transfer performance of SOTA deep learning computer vision models for UOAIS. These results highlight the potential of SynTable for addressing the challenge of annotating amodal instance segmentation masks.